OpenStudy (kalijah13):

Lizette’s math test grades are 70, 72, and 68. If she scores 100 on her next test, what will happen to the RANGE of her scores? A) It will double. B) It will stay the same. C) It will increase 6 times. D) It will increase 8 times.

OpenStudy (kalijah13):

PLZ HELP I WILL MEDAL

OpenStudy (caozeyuan):

range is biggest minus smallest

OpenStudy (kalijah13):

that is correct

OpenStudy (caozeyuan):

range before =72-68=4

OpenStudy (kalijah13):

ohhhhhh okay

OpenStudy (caozeyuan):

range aftter is 100-68=32

OpenStudy (kalijah13):

so the answer would be... A????

OpenStudy (caozeyuan):

what is 32/4

OpenStudy (kalijah13):

OOOOOHHHHHHH

OpenStudy (kalijah13):

MY GOD IM SO DUMB

OpenStudy (kalijah13):

8

OpenStudy (caozeyuan):

its okay, I feel dumb all the time

OpenStudy (caozeyuan):

so answer is ....

OpenStudy (kalijah13):

But you helped me and i am grateful thank you.

OpenStudy (caozeyuan):

Actually I am in liz's situatio right now, hope I can get 100 next time